# GP-PCS: One-shot Feature-Preserving Point Cloud Simplification with Gaussian Processes on Riemannian Manifolds

### v3

@@ Abstract
The processing, storage and transmission of large-scale point clouds is an ongoing challenge in the computer vision community which hinders progress in the application of 3D models to real-world settings, such as autonomous driving, virtual reality and remote sensing. We propose a novel, one-shot point cloud simplification method which preserves both the salient structural features and the overall shape of a point cloud without any prior surface reconstruction step. Our method employs Gaussian processes suitable for functions defined on Riemannian manifolds, allowing us to model the surface variation function across any given point cloud. A simplified version of the original cloud is obtained by sequentially selecting points using a greedy sparsification scheme. The selection criterion used for this scheme ensures that the simplified cloud best represents the surface variation of the original point cloud. We evaluate our method on several benchmark and self-acquired point clouds, compare it to a range of existing methods, demonstrate its application in downstream tasks of registration and surface reconstruction, and show that our method is competitive both in terms of empirical performance and computational efficiency.
$\dagger$\$$\dagger$\$footnotetext: Equal contribution.
@@ 1 Introduction
Figure 1: Point cloud simplification methods typically fail to strike a balance between preserving sharp features and maintaining the overall structure of the original cloud. Our approach circumvents this trade-off by achieving both targets, as is evident from the simplified versions of the Stanford Bunny [20] obtained using the proposed technique and two pre-existing methods; Hierarchical Clustering (HC) and Weighted Locally Optimal Projection (WLOP).
@@
In this work, we propose a novel, one-shot, feature-preserving simplification method using GPs with kernels defined on Riemannian manifolds. Using a greedy algorithm for GP sparsification, we iteratively construct a simplified representation of a point cloud without the need for any prior surface reconstruction or training on large point cloud datasets. We experiment on several point clouds, compare with several techniques and demonstrate competitive results both empirically and in terms of computational efficiency. Qualitatively, as shown in Fig. 1, our method effectively preserves visual features whilst providing a sufficiently dense coverage of the domain of the original cloud.
Outline of the paper: Section 2 briefly reviews a number of existing point cloud simplification techniques which are relevant to our work. Section 3 provides background details regarding the computation of surface variation, GPs with kernels defined on non-Euclidean domains and a greedy subset-of-data scheme for GP inference. Section 4 outlines the proposed GP-based point cloud simplification algorithm. Section 5, in combination with the supplementary material, includes an empirical evaluation of our method on various benchmark and self-acquired point clouds, with comparisons to competing simplification techniques, along with applications to some downstream tasks. Finally, Section 6 summarises our contributions and provides a brief discussion of the scope for future work.
@@ 2 Related work
Another widely-used technique is Weighted Locally Optimal Projection (WLOP) proposed by Huang et al. [14]. In this work, the authors modified the existing parameterization-free denoising simplification scheme termed Locally Optimal Projection (LOP) [22], which is unsuitable for non-uniformly distributed point clouds. WLOP overcomes this limitation by incorporating locally adaptive density weights into LOP. Although WLOP results in an evenly distributed simplified cloud, it still lacks sensitivity towards salient geometric features which will also become apparent in Section 5. Recently, Potamias et al. [28] have proposed a graph neural network-based learnable simplification technique which uses a modified variant of Chamfer distance in order to backpropagate errors. Their method can simplify point clouds in real-time but involves a computationally intensive training process using large point cloud datasets such as TOSCA [6]. Moreover, their model’s efficiency is limited to simplifying point clouds which are structurally similar to the learned data, as inherently neural networks struggle to generalize outside of the domain of the training data.
Approximate Intrinsic Voxel Structure for Point Cloud Simplification (AIVS), introduced by Lv et al. [24], combines global voxel structure and local farthest point sampling to generate simplification demand-specific clouds which can be either isotropic, curvature-sensitive or have sharp edge preservation. As with HC however, AIVS requires manual tuning of user-specified parameters in order to obtain optimal results. Additionally, even in parallel computation mode, AIVS is quite costly in terms of computational runtime. Potamias et al. and Lv et al. do not provide open-source implementations of their curvature-sensitive simplification techniques, which poses a challenge for reproducibility and benchmarking. Qi et al. [29] introduced PC-Simp, a method which aims to produce uniformly-dense and feature-sensitive simplified clouds, leveraging ideas from graph signal processing. This uniformity depends on a weight parameter which as with HC and AIVS, is user-specified. Alongside simplification, they also apply their technique to point cloud registration. However, in practice PC-Simp is unreliable for complex-surfaced point clouds as it fails to provide a high degree of feature-preservation, regardless of the weight parameter chosen. Additionally, as discussed later in Section 5, the runtime of this technique is considerably longer than any other method tested.
@@ 3.1 Surface variation
Consider an unstructured dense point cloud $P=\{\boldsymbol{p_{1}},\boldsymbol{p_{2}},...,\boldsymbol{p_{N}}\}$ of size $N$ existing in 3D Euclidean space, $\mathbb{R}^{3}$. We can generate the local neighbourhood $N_{\boldsymbol{p_{i}}}$ of each point $\boldsymbol{p_{i}}$ in $P$ by two different methods. Firstly, we can gather all of the points within a certain Euclidean distance $r$ from $\boldsymbol{p_{i}}$; this approach is referred to as radius search. Alternatively, we can gather all of the k-nearest Euclidean neighbours of $\boldsymbol{p_{i}}$, which is referred to as KNN search. The choice of this scale-factor ($r$ or $k$) not only depends on the size and density of a point cloud but also on the desired level of detail for a given application. These aspects make the task of automatic estimation of the neighbourhood of a point in a cloud an important, yet challenging one [31]. In this work, we implement the approach taken by the CloudCompare software package, where this process is automated by first calculating an approximate surface per point from the bounding box volume. This estimated value, along with a user-defined approximate neighbour number, is used to estimate a radius $r$, which is then used to perform radius search for each point. In our method, we have fixed this approximate neighbour number to $25$ as it provides good empirical performance across a wide variety of point clouds.
@@ 3.2 Gaussian processes on Riemannian manifolds
Inference: Using Bayes’ Rule, we can condition our GP on the training data and derive closed form expressions for the posterior mean and covariance:
@@
Kernels on manifolds: Many different kernel functions for GPs exist, and choosing a kernel is in itself a model selection problem as some kernels are more suited to modeling certain types of data. However, one characteristic which many kernels share is that they are defined using Euclidean distance. This presents an issue should we wish to use a GP to model variation in a quantity over a non-Euclidean space. Borovitskiy et al. [5] proposed a solution to this problem in the form of an extension to the Matérn kernel, which allows for modeling of functions whose domains are compact Riemannian manifolds. The approach proposed by the authors involves two stages. Firstly, numerical estimation of the eigenvalues $\lambda_{n}$ and eigenfunctions $f_{n}$ corresponding to the Laplace-Beltrami operator of the given manifold is performed. Secondly, for a manifold of dimensionality $d$, the kernel is approximated using a finite truncation of:
@@ 4 Point cloud simplification with Riemannian Gaussian processes
Table 1: Empirical results and total runtimes (time taken by surface variation computation and simplification) for all tested simplification methods and point clouds. We report the maximum and mean Hausdorff distances between the original meshes, and the meshes reconstructed from the simplified point clouds. Also reported is the average surface variation over each simplified point cloud. Best results are boldfaced, whilst second-best are underlined.
@@
5 Empirical evaluation
In this section, we extensively evaluate the proposed simplification method using various point cloud datasets and processing techniques. First, we compare our simplification technique both quantitively and qualitatively using benchmark object level point clouds as given in subsection 5.1. Second, in subsection 5.2 we extend the use-case of our algorithm as a time and memory efficient pre-processing step for the downstream task of point cloud registration. Moreover, we provide some experiments on scene level and self-acquired point clouds in the supplementary material (Section 2).
5.1 Benchmark object level point clouds
Evaluation criteria: In order to evaluate the performance of our method in comparison to other simplification techniques, we firstly use each simplified point cloud obtained from three object level point clouds to form simplified meshes, using screened Poisson surface reconstruction [17]. We can then compute the reconstruction errors between the original meshes, and the reconstructed meshes formed from our simplified clouds. Specifically, we choose to evaluate the mean and maximum Hausdorff distance [8]. Evaluating the error associated with mesh reconstruction is effective at quantifying the ability of each method to preserve features from the original cloud, as accurate reconstruction of a mesh from a simplified point cloud requires that a high density of points be placed in the vicinity of finer details within the cloud. The MeshLab software was used to reconstruct all surfaces and compute the Hausdorff distances. Also, given that one of our primary aims is to preserve sharp features within each point cloud, we also report the average surface variation over each simplified point cloud. The surface variation at each point is computed using the approach described in Section 3.1.
Baselines: We use the aforementioned evaluation procedure to compare our method (denoted GP) empirically to a number of competing simplification techniques discussed in Section 2. We compare our approach to PC-Simp, AIVS, HC and WLOP, with the latter two approaches implemented using the CGAL library. For the HC method, the size and variation parameters discussed in Section 2 were manually tuned to obtain approximate desired simplified sizes. Also, as noted in Section 2, we use the non-curvature aware version of the AIVS algorithm, as there is no available open-source implementation of the curvature-aware variant.
Experimental details: We evaluate our proposed method and the aforementioned baselines on three complex object-level point clouds from the Stanford 3D Scanning Repository [20], namely Armadillo ($N=1,72,974$), Dragon ($N=4,37,645$) and Lucy ($N=1,40,27,872$). Let the simplification ratio be defined as $\alpha=M/N$. In this work we focus on the challenging regime where we wish to significantly reduce the size of the cloud, such that $\alpha<<1$. It is in this regime that feature-preserving techniques such as ours become particularly important, as we do not have a large number of points to select, thus we must efficiently select points which allow us to capture the salient features of the original cloud. We chose $\alpha$ for each cloud by finding the minimum $\alpha$ at which all evaluated techniques were capable of forming simplified clouds from which meshes visually comparable to the original meshes could be generated [20]. This value varies depending on the surface complexity of each cloud, thus for Armadillo, Dragon and Lucy we chose $\alpha=0.05$, $0.03$ and $0.002$ respectively. Additionally, we also visually evaluate the point cloud simplification results of all aforementioned techniques on a noisy Armadillo from the PCPNet dataset [12], with $\alpha=0.05$. This corresponds to the original Armadillo model surface sampled $10^{5}$ times ($N=1,00,000$), with Gaussian noise (of standard deviation $\sigma=2.5\times 10^{-3}\times d$, where $d$ is bounding box diagonal length) added to every point position. Our code, which will be open sourced upon publication, is included in the supplementary material alongside further experimental details (Section 1).
Discussion: From the results presented in Table 1, it is clear that our proposed method is capable of comparable empirical performance to many of the existing methods for simplifying point clouds. The GP-based approach outperforms the AIVS baseline across all experiments and metrics, and outperforms the PC-Simp baseline on all but the mean Hausdorff distance for the Armadillo experiment. Moreover, our algorithm also runs considerably faster than both of these approaches. Note that due to the scale of Lucy, we were unable to evaluate PC-Simp on this cloud as it was taking more than two hours to run.
@@ 5.1 Benchmark object level point clouds
HC is the only baseline with a shorter runtime than our method, and obtains maximum Hausdorff distances comparable to those obtained by our approach. However, as discussed in Section 2, tuning the user-specified HC parameters make striking a balance between feature preservation and retaining a sufficient density of points across the cloud relatively challenging. Moreover, there is no control over the size of the simplified cloud, as discussed by the authors [26] and in subsequent work [24]. We tuned this baseline to attempt to balance this trade-off, and whilst the HC-simplified clouds shown in Figures 2 and 3 here, and Figure 3 of the supplementary material, do have clearly preserved features (an observation supported by the high mean surface variation across all clouds), the density of points away from these areas is very low. This leads to inferior mesh reconstructions compared to our approach, as evidenced by the fact that we obtain superior mean Hausdorff distance compared to HC across all three clouds.
@@
Overall, these results show that our approach provides a computationally efficient option for performing point cloud simplification in settings where the user wishes to strike a balance between preserving high fidelity around sharp features in the cloud, and ensuring that the simplified cloud covers the manifold defined by the original cloud with a sufficient density of points. This is important for generating reconstructions which resemble the original meshes, as is evident from visual inspection of the reconstruction results in Figure 2 here and Figure 3 of the supplementary material. In terms of surface reconstruction, our method clearly outperforms all of the other techniques for the Dragon (compare the tail, teeth, horns and the face detailing for all methods and additionally the curved body for HC) and the Armadillo (compare the ears, hands and feet across all the methods) and gives competitive results for Lucy, shown in Figure 2 of the supplementary material. Again, visual inspection of the simplification results for the noisy armadillo in Figure 3 demonstrates the balanced feature-sensitivity of our method in comparison to others.

### v4

@@ Abstract
The processing, storage and transmission of large-scale point clouds is an ongoing challenge in the computer vision community which hinders progress in the application of 3D models to real-world settings, such as autonomous driving, virtual reality and remote sensing. We propose a novel, one-shot point cloud simplification method which preserves both the salient structural features and the overall shape of a point cloud without any prior surface reconstruction step. Our method employs Gaussian processes suitable for functions defined on Riemannian manifolds, allowing us to model the surface variation function across any given point cloud. A simplified version of the original cloud is obtained by sequentially selecting points using a greedy sparsification scheme. The selection criterion used for this scheme ensures that the simplified cloud best represents the surface variation of the original point cloud. We evaluate our method on several benchmark and self-acquired point clouds, compare it to a range of existing methods, demonstrate its application in downstream tasks of registration and surface reconstruction, and show that our method is competitive both in terms of empirical performance and computational efficiency. The code is available at https://github.com/stutipathak5/gps-for-point-clouds.
Keywords:
Point clouds Simplification Gaussian processes Riemannian manifolds
@@ 1 Introduction
Figure 1: Point cloud simplification methods typically fail to strike a balance between preserving sharp features and maintaining the overall structure of the original cloud. Our approach circumvents this trade-off by achieving both targets, as is evident from the simplified versions of the Stanford Bunny [20] obtained using the proposed technique and three pre-existing methods; Hierarchical Clustering (HC) [26], Weighted Locally Optimal Projection (WLOP) [14], and Potamias et al. [28] simplification.
In this work, we propose a novel, one-shot, feature-preserving simplification method using GPs with kernels defined on Riemannian manifolds. Using a greedy algorithm for GP sparsification, we iteratively construct a simplified representation of a point cloud without the need for any prior surface reconstruction or training on large point cloud datasets. We experiment on several point clouds, compare with several techniques and demonstrate competitive results both empirically and in terms of computational efficiency. Qualitatively, as shown in Figure 1, our method effectively preserves visual features whilst providing a sufficiently dense coverage of the domain of the original cloud.
1.0.1 Outline of the paper:
Section 2 briefly reviews a number of existing point cloud simplification techniques which are relevant to our work. Section 3 provides background details regarding the computation of surface variation, GPs with kernels defined on non-Euclidean domains and a greedy subset-of-data scheme for GP inference. Section 4 outlines the proposed GP-based point cloud simplification algorithm. Section 5, in combination with the supplementary material, includes an empirical evaluation of our method on various benchmark and self-acquired point clouds, with comparisons to competing simplification techniques, along with applications to some downstream tasks and ablation studies. Finally, Section 6 summarises our contributions and provides a brief discussion of the scope for future work.
@@ 2 Related work
Another widely-used technique is Weighted Locally Optimal Projection (WLOP) proposed by Huang et al. [14]. In this work, the authors modified the existing parameterization-free denoising simplification scheme termed Locally Optimal Projection (LOP) [22], which is unsuitable for non-uniformly distributed point clouds. WLOP overcomes this limitation by incorporating locally adaptive density weights into LOP. Although WLOP results in an evenly distributed simplified cloud, it still lacks sensitivity towards salient geometric features which will also become apparent in Section 5. Recently, Potamias et al. [28] have proposed a graph neural network-based learnable simplification technique which uses a modified variant of Chamfer distance in order to backpropagate errors. Their method can simplify point clouds in real-time but involves a computationally intensive training process using large point cloud datasets such as TOSCA [6]. Moreover, their model’s efficiency is limited to simplifying point clouds which are structurally similar to the learned data, as inherently neural networks struggle to generalize outside of the domain of the training data. Even more recent work from Wu et al. [33], named APES, proposes an edge-sampling method which claims to capture the salient points within a point cloud using an attention mechanism. As shown in their paper, this technique generally provides good results for some point cloud tasks. However, as discussed by the authors themselves, the edge-enhancing nature of their method hinders upsampling operations, which can lead to poor reconstruction and segmentation results later.
Approximate Intrinsic Voxel Structure for Point Cloud Simplification (AIVS), introduced by Lv et al. [24], combines global voxel structure and local farthest point sampling to generate simplification demand-specific clouds which can be either isotropic, curvature-sensitive or have sharp edge preservation. As with HC however, AIVS requires manual tuning of user-specified parameters in order to obtain optimal results. Additionally, even in parallel computation mode, AIVS is quite costly in terms of computational runtime. Potamias et al. and Lv et al. do not provide open-source implementations of their curvature-sensitive simplification techniques, which poses a challenge for reproducibility and benchmarking. However, we thank the authors of Potamias et al. for directly providing some simplified point clouds; their results are included later in this paper. Qi et al. [29] introduced PC-Simp, a method which aims to produce uniformly-dense and feature-sensitive simplified clouds, leveraging ideas from graph signal processing. This uniformity depends on a weight parameter which as with HC and AIVS, is user-specified. Alongside simplification, they also apply their technique to point cloud registration. However, in practice PC-Simp is unreliable for complex-surfaced point clouds as it fails to provide a high degree of feature-preservation, regardless of the weight parameter chosen. Additionally, as discussed later in Section 5, the runtime of this technique is considerably longer than any other method tested.
@@ 3.1 Surface variation
Consider an unstructured dense point cloud $P=\{\boldsymbol{p_{1}},\boldsymbol{p_{2}},...,\boldsymbol{p_{N}}\}$ of size $N$ existing in 3D Euclidean space, $\mathbb{R}^{3}$. We can generate the local neighbourhood $N_{\boldsymbol{p_{i}}}$ of each point $\boldsymbol{p_{i}}$ in $P$ by two different methods. Firstly, we can gather all of the points within a certain Euclidean distance $r$ from $\boldsymbol{p_{i}}$; this approach is referred to as radius search. Alternatively, we can gather all of the k-nearest Euclidean neighbours of $\boldsymbol{p_{i}}$, which is referred to as KNN search. The choice of this scale-factor ($r$ or $k$) not only depends on the size and density of a point cloud but also on the desired level of detail for a given application. These aspects make the task of automatic estimation of the neighbourhood of a point in a cloud an important, yet challenging one [31]. In this work, we implement the approach taken by the CloudCompare software package, where this process is automated by first calculating an approximate surface per point from the bounding box volume. This estimated value, along with a user-defined approximate neighbour number, is used to estimate a radius $r$, which is then used to perform radius search for each point. In our method, we have fixed this approximate neighbour number to $25$ as it provides good empirical performance across a wide variety of point clouds. However, we provide ablation studies over a range of neighbourhood sizes in the supplementary material.
@@ 3.2 Gaussian processes on Riemannian manifolds
3.2.1 Inference:
Using Bayes’ Rule, we can condition our GP on the training data and derive closed form expressions for the posterior mean and covariance:
@@ 3.2.1 Inference:
3.2.2 Kernels on manifolds:
Many different kernel functions for GPs exist, and choosing a kernel is in itself a model selection problem as some kernels are more suited to modeling certain types of data. However, one characteristic which many kernels share is that they are defined using Euclidean distance. This presents an issue should we wish to use a GP to model variation in a quantity over a non-Euclidean space. Borovitskiy et al. [5] proposed a solution to this problem in the form of an extension to the Matérn kernel, which allows for modeling of functions whose domains are compact Riemannian manifolds. The approach proposed by the authors involves two stages. Firstly, numerical estimation of the eigenvalues $\lambda_{n}$ and eigenfunctions $f_{n}$ corresponding to the Laplace-Beltrami operator of the given manifold is performed. Secondly, for a manifold of dimensionality $d$, the kernel is approximated using a finite truncation of:
@@ 4 Point cloud simplification with Riemannian Gaussian processes
5 Empirical evaluation
In this section, we extensively evaluate the proposed simplification method using various point cloud datasets and processing techniques. First, we compare our simplification technique both quantitively and qualitatively using benchmark object level point clouds as given in subsection 5.1. Second, in subsection 5.2 we extend the use-case of our algorithm as a time and memory efficient pre-processing step for the downstream task of point cloud registration. Moreover, we provide some experiments on scene level and self-acquired point clouds along with ablation studies in the supplementary material (Section 2).
5.1 Benchmark object level point clouds
5.1.1 Evaluation criteria:
In order to evaluate the performance of our method in comparison to other simplification techniques, we firstly use each simplified point cloud obtained from three object level point clouds to form simplified meshes, using screened Poisson surface reconstruction [17]. We can then compute the reconstruction errors between the original meshes, and the reconstructed meshes formed from our simplified clouds. Specifically, we choose to evaluate the mean and maximum Hausdorff distance [8]. Evaluating the error associated with mesh reconstruction is effective at quantifying the ability of each method to preserve features from the original cloud, as accurate reconstruction of a mesh from a simplified point cloud requires that a high density of points be placed in the vicinity of finer details within the cloud. The MeshLab software was used to reconstruct all surfaces and compute the Hausdorff distances. Also, given that one of our primary aims is to preserve sharp features within each point cloud, we also report the average surface variation over each simplified point cloud. The surface variation at each point is computed using the approach described in Section 3.1.
5.1.2 Baselines:
We use the aforementioned evaluation procedure to compare our method (denoted GP) empirically to a number of competing simplification techniques discussed in Section 2. We compare our approach to PC-Simp, AIVS, Potamias et al., HC and WLOP, with the latter two approaches implemented using the CGAL library. Additionally, we provide a visual comparison of our simplification method with APES. For the HC method, the size and variation parameters discussed in Section 2 were manually tuned to obtain approximate desired simplified sizes. Also, as noted in Section 2, we use the non-curvature aware version of the AIVS algorithm, as there is no available open-source implementation of the curvature-aware variant.
5.1.3 Experimental details:
We evaluate our proposed method and the aforementioned baselines on three complex object-level point clouds from the Stanford 3D Scanning Repository [20], namely Armadillo ($N=1,72,974$), Dragon ($N=4,37,645$) and Lucy ($N=1,40,27,872$). Let the simplification ratio be defined as $\alpha=M/N$. In this work we focus on the challenging regime where we wish to significantly reduce the size of the cloud, such that $\alpha<<1$. It is in this regime that feature-preserving techniques such as ours become particularly important, as we do not have a large number of points to select, thus we must efficiently select points which allow us to capture the salient features of the original cloud. We chose $\alpha$ for each cloud by finding the minimum $\alpha$ at which all evaluated techniques were capable of forming simplified clouds from which meshes visually comparable to the original meshes could be generated [20]. This value varies depending on the surface complexity of each cloud, thus for Armadillo, Dragon and Lucy we chose $\alpha=0.05$, $0.03$ and $0.002$ respectively. Additionally, we also visually evaluate the point cloud simplification results of all aforementioned techniques on a noisy Armadillo from the PCPNet dataset [12], with $\alpha=0.05$. This corresponds to the original Armadillo model surface sampled $10^{5}$ times ($N=1,00,000$), with Gaussian noise (of standard deviation $\sigma=2.5\times 10^{-3}\times d$, where $d$ is bounding box diagonal length) added to every point position. We also perform the same evaluation for three objects, an airplane, a glass and a toilet ($N=2,048$ for all three) from ModelNet40 dataset [34] to compare our method with APES.
Figure 4: Simplification results of an airplane, a glass and a toilet point cloud for simplification ratios $\alpha=0.25$ and $0.5$ using APES and GP-based simplification.
5.1.4 Discussion:
From the results presented in Table 1, it is clear that our proposed method is capable of comparable empirical performance to many of the existing methods for simplifying point clouds. The GP-based approach outperforms the AIVS baseline across all experiments and metrics, and outperforms the PC-Simp baseline on all but the mean Hausdorff distance for the Armadillo experiment. Moreover, our algorithm also runs considerably faster than both of these approaches. Note that due to the scale of Lucy, we were unable to evaluate PC-Simp on this cloud as it was taking more than two hours to run.
@@ 5.1.4 Discussion:
Table 1: Empirical results and total runtimes (time taken by surface variation computation and simplification) for all tested simplification methods and point clouds. We report the maximum and mean Hausdorff distances between the original meshes, and the meshes reconstructed from the simplified point clouds. Also reported is the average surface variation over each simplified point cloud. Best, second-best and third-best results are in red, green and blue respectively. It is worth mentioning that as per the evaluated metrics, our algorithm mostly stays within the top three methods.
@@
Potamias et al. 0.215 0.000599 4.28 3.47 0.00933 190.00 0.0478 0.0650 0.0511 0.00060 0.00070 0.00212
@@
HC and Potamias et al. are the only baselines with shorter runtimes than our method, and obtain maximum Hausdorff distances comparable to those obtained by our approach. However, as discussed in Section 2, tuning the user-specified HC parameters make striking a balance between feature preservation and retaining a sufficient density of points across the cloud relatively challenging. Moreover, there is no control over the size of the simplified cloud, as discussed by the authors [26] and in subsequent work [24]. We tuned this baseline to attempt to balance this trade-off, and whilst the HC-simplified clouds shown in Figures 2 and 3 here, and Figure 3 of the supplementary material, do have clearly preserved features (an observation supported by the high mean surface variation across all clouds), the density of points away from these areas is very low. This leads to inferior mesh reconstructions compared to our approach, as evidenced by the fact that we obtain superior mean Hausdorff distance compared to HC across all three clouds.
Since results and inference times for the Potamias et al. approach were provided by the author of the paper, we do not have knowledge of the exact details of their experimental setup, especially the time required in hours to train the model. As mentioned in Section 2, their learning-based approach demands huge datasets to train on, which not only increases the computational requirements but also limits their method’s generalizability. When compared to our method quantitatively, our method generally gives superior results, except for two of the nine error metric values. This is supported by the quality of their simplified point clouds and the corresponding reconstructions shown in Figures 2 and 3 here, and in Figures 2, 3 and 5 of the supplementary material. Although their method performs best in the case of Lucy’s maximum Hausdorff distance, in reality their simplified cloud gives arguably the poorest qualitative reconstruction result amongst all of the other baselines. As expected, the inference time of their approach is the lowest of all the baselines, because of their neural network-based approach, which involves pre-training.
@@
Overall, these results show that our approach provides a computationally efficient option for performing point cloud simplification in settings where the user wishes to strike a balance between preserving high fidelity around sharp features in the cloud, and ensuring that the simplified cloud covers the manifold defined by the original cloud with a sufficient density of points. This is important for generating reconstructions which resemble the original meshes, as is evident from visual inspection of the reconstruction results in Figure 2 here and Figure 3 of the supplementary material. In terms of surface reconstruction, our method clearly outperforms all of the other techniques for the Dragon (compare the tail, teeth, horns and the face detailing for all methods and additionally the curved body for HC) and the Armadillo (compare the ears, hands and feet across all the methods) and gives competitive results for Lucy, shown in Figure 2 of the supplementary material. We highlight once again the poor surface reconstructions resulting from the Potamias et al. simplified clouds, compared to those obtained using all of the other baselines. Again, visual inspection of the simplification results for the noisy Armadillo in Figure 3 demonstrates the balanced feature-sensitivity of our method in comparison to others. We experiment with more noise levels in the supplementary material (Section 2). Finally, from Figure 4 we can see how the edge-sampling-based APES simplified clouds have several missing portions including object edges, whereas our method enhances the salient features and captures the overall object structure simultaneously. We do not provide corresponding surface reconstructions and hence quantitative results for this baseline because their low simplified point cloud sizes ($N=1,024$ and $512$) and aforementioned missing areas will always result in open meshes.